4 value: 0.58 points For each of the following independent cases A-D, compute the missing values in the table below Case Materias Direct Labort Prime Cost Convession Direct Materials Case Overhead Prime Cost Cost 900$ 1,300 $ 2,000 1,325 1,500 2,650 2,880 1,600 400 700 750 2,0009 value: 0.58 points For each of the following independent cases (A-E), compute the missing values in the table Total Prime Cost Conversion Cost Direct Materials Direct Manufacturing Labor Manufacturing Cost ase Overhead $2,130$1,060 $ 3,600 6,950 11,670 7,970 2,160 6,880 1,520 3,240 9,490 2,020 2,920 5,940 11,580 20,610 3,390please answer both
Answer #1

Requirement 1

Case

Prime cost

Conversion cost

Direct material

Direct labor

Manufacturing overhead

Total Manufacturing cost

A

$   3,190.00

$    4,660.00

$ 2,130.00

$ 1,060.00

$    3,600.00

$   6,790.00

B

$   6,950.00

$ 11,670.00

$ 2,160.00

$ 4,790.00

$    6,880.00

$ 13,830.00

C

$   4,760.00

$    7,970.00

$ 1,520.00

$ 3,240.00

$    9,490.00

$   9,490.00

D

$   3,020.00

$    4,940.00

$ 1,000.00

$ 2,020.00

$    2,920.00

$   5,940.00

E

$ 11,580.00

$ 20,610.00

$ 3,390.00

$ 8,190.00

$ 12,420.00

$ 24,000.00

Working

Case

Prime cost

Conversion cost

Direct material

Direct labor

Manufacturing overhead

Total Manufacturing cost

A

=2130+1060

=1060+3600

2130

1060

3600

=3600+1060+2130

B

6950

11670

2160

=11670-6880

6880

=6880+4790+2160

C

=1520+3240

7970

1520

3240

=1520+7970

9490

D

=1000+2020

=2020+2920

=5940-4940

2020

2920

5940

E

11580

20610

3390

=11580-3390

=20610-8190

=12420+8190+3390

Requirement 2

Case

Direct material

Direct labor

Manufacturing Overheads

Prime cost

Conversion cost

A

$    900.00

$    1,300.00

$ 2,000.00

$ 2,200.00

$ 3,300.00

B

$     400.00

$    2,250.00

$ 1,325.00

$ 2,650.00

$ 3,575.00

C

$ 2,180.00

$       700.00

$ 1,500.00

$ 2,880.00

$ 2,200.00

D

$     850.00

$       750.00

$ 1,250.00

$ 1,600.00

$ 2,000.00

Working

Case

Direct material

Direct labor

Manufacturing Overheads

Prime cost

Conversion cost

A

900

1300

2000

=900+1300

=1300+2000

B

400

=2650-400

1325

2650

=2250+1325

C

=2880-700

700

1500

2880

=700+1500

D

=1600-750

750

=2000-750

1600

2000

Conversion cost = Direct labor and manufacturing overheads

Prime cost= Direct material +Direct labor